### v2

# The dark side of the universe may be more harmonic than we thought

Abstract
The standard paradigm of cosmology assumes two distinct dark components, namely the dark energy driving the late-universe acceleration and the dark matter that is responsible for the structure formation. However, the necessity of splitting the dark-side world into two sectors has not been experimentally or theoretically proven. It is shown in Wang et al. 2024 that cosmology with one unified dark fluid can also explain the cosmic microwave background (CMB) and late-universe data, with the fitting quality not much worse than the standard Lambda cold dark matter ($\Lambda$CDM) model. The present work aims to provide a clearer physical interpretation of the Wang et al. 2024 results. We show that the unified dark fluid model can produce primary CMB temperature and polarization power spectra that are very close to the $\Lambda$CDM prediction (relative difference $\lesssim 10^{-4}$). The model can also mimic the $\Lambda$CDM background expansion history and linear growth factor on sub-horizon scales with percent-level accuracy. With better physical understanding of the model, we make precision tests and find a minor error in the Boltzmann code used in Wang et al. 2024. We correct the error and update the model comparison between $\Lambda$CDM and the unified dark fluid model.
preprint: APS/123-QED
I Introduction
Our universe contains approximately 5% baryonic matter and 95% dark components which are commonly considered as dark matter and dark energy [1]. Dark matter plays an important role in the formation of large scale structures, while dark energy drives the accelerated expansion of the universe [2, 3]. In the standard Lambda cold dark matter ($\Lambda$CDM) model, dark energy is interpreted as the cosmological constant ($\Lambda$) or equivalently the vacuum energy. The cosmological constant interpretation of dark energy has a fine-tuning problem, which questions the smallness of vacuum energy density [4], and a coincidence problem, which asks why the vacuum energy density is the same order of magnitude as the matter density today [5]. The fine-tuning and coincidence problems also apply to many alternative models of dark energy [6, 7].
The coincidence between the densities of dark matter and baryonic matter is usually considered to be less problematic, as baryons and dark matter may have a similar origin in the early universe. Thus, the coincidence problem of dark energy could be naturally resolved if we unify dark energy and dark matter into one single component that shares a common origin with baryonic matter. To explain the cosmological data, the unified dark component should behave like pressure-less dust in the early (redshift $z\gg 1$) universe and should have negative pressure in the late ($z\lesssim 1$) universe. If the dust-to-$\Lambda$ transition could be triggered by the inhomogeneity of the unified dark component itself, or by its coupling to neutrinos which becomes non-relativistic in the late-universe, the fine-tuning problem would also be resolved.
Beyond enduring conceptual difficulties, the vacuum-energy paradigm for dark energy has been further called into question by recent advances in observational cosmology. Recent multi-probe analyses incorporating the Dark Energy Spectroscopic Instrument (DESI) baryon acoustic oscillations (BAO) measurements, Type Ia supernovae (SNe) luminosity distances, and cosmic microwave background (CMB) power spectra provide strong empirical support for phantom-crossing dark energy scenarios [8, 9]. Theoretically, constructing such phantom-divide crossing behavior within a single-component framework poses significant challenges, as canonical scalar field implementations typically develop quantum instabilities when approaching the phantom threshold (press-to-density ratio less than $-1$) due to effective negative kinetic energy terms. Notably, the effective equation of state (total pressure divided by total density) for the combined dark sector (dark matter plus dark energy) in the DESI BAO+SNe+CMB best-fit model remains above $-1$ across cosmic history. This again motivates the unification paradigm wherein dark matter and dark energy emerge as different manifestations of a unified dark component, thereby naturally circumventing the phantom-crossing problem.
While it is difficult to formulate a fundamental theory to implement all the aforementioned ideas, it is possible to construct an effective action or to build a phenomenological model with fluid approximation. Examples include Chaplygin gas and its many variations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], scalar field with non-canonical kinetic energy [25, 26, 27, 28, 29, 30, 31, 32], modified gravity theories [33, 34, 35, 36, 37, 38, 39, 40, 41], quark bag model [42], Bose-Einstein condensate [43], polytropic dark matter [44], and other fluid models [45, 46, 47, 48, 49]. Although some of the models have difficulties to predict cosmological perturbations that fit the current data [50, 51, 52, 53, 54], it has been shown numerically that a unified dark fluid with negligible anisotropic stress and zero sound speed in general can make $\Lambda$CDM-like predictions at background and linear-perturbations levels [55, 49].
In the PAge-like unified dark fluid (PUDF) model that was proposed in Wang et al. [49], the unified dark component is assumed to be a fluid with a smooth background evolution parameterized by the PAge approximation [56, 57]. The PAge approximation is based on two assumptions, that the dark component(s) behave like dust at high-redshift, and that the dimensionless combination $Ht$, where $H$ is the Hubble parameter and $t$ is the age of the universe, is a slowly varying smooth function of $t$. The minimal PUDF contains seven cosmological parameters, with the standard $\Omega_{c}h^{2}$ (CDM density) replaced by the PAge parameters $p_{\rm age}$ ($\sim$ age of the universe) and $\eta$ (deviation from Einstein-de Sitter universe). By modifying the Boltzmann code CLASS [58], Wang et al. [49] computed the linear perturbations in PUDF and found that PUDF can give predictions similar to those of $\Lambda$CDM. Further analysis of Bayesian evidence shows that $\Lambda$CDM is favored over PUDF by the current cosmological data including CMB, BAO, SNe, and cosmic chronometers (CC) [49].
The results found in Wang et al. [49], however, lack a clear physical interpretation. It is unclear to what extent PUDF can mimic $\Lambda$CDM at the background and linear-perturbation levels. Neither do we know what key difference between PUDF and $\Lambda$CDM has led to the slightly different $\chi^{2}$ fits to the data. Similar problems exist for the earlier work Davari et al. [55] with a polynomial-based parameterization. This work then aims to improve the theoretical understanding of the similarities and nuances between PUDF and $\Lambda$CDM, and to come up with some quantitative predictions that can be used to test the numerical accuracy of the Boltzmann code. While the theoretical exploration is done in Section II, we revisit the Bayesian parameter inference and update some of the results in Section III. Section IV summarizes and concludes.
Throughout the paper we work with the spatially flat background metric $\mathrm{d}s^{2}=\mathrm{d}t^{2}-a(t)^{2}\mathrm{d}\mathbf{x}^{2}$, where the scale factor $a(t)$ is related to the cosmological redshift $z$ via $a=\frac{1}{1+z}$. The Hubble parameter is defined as $H(t)=\frac{\dot{a}}{a}$, where a dot denotes derivative with respect to the background time $t$. We use a subscript 00 to denote quantities at redshift zero. For example, the Hubble constant $H_{0}$ is the Hubble parameter at redshift zero, often written as $100h\,\mathrm{km\cdot s^{-1}Mpc^{-1}}$. The critical density is defined as $\rho_{\rm crit}=\frac{3H_{0}^{2}}{8\pi G}$, where $G$ is Newton’s gravitational constant. We use subscripts $b$, $c$, $d$, $\nu$, $\gamma$, $\Lambda$ for baryon, cold dark matter, unified dark fluid, neutrinos, photons and vacuum energy, respectively. For a component $X=b,c,d,\nu,\gamma,\Lambda$, the abundance parameter $\Omega_{X}$ is defined as the ratio between its current background density $\rho_{X0}$ and the critical density $\rho_{\rm crit}$. For parameter inference, unless otherwise specified, we assume flat priors on the logarithm amplitude of primordial scalar perturbations $\ln(10^{10}A_{s})$, the tilt of primordial scalar perturbations $n_{s}$, the reionization optical depth $\tau_{\rm re}$, the angular extension of the sound horizon at recombination $\theta_{*}$, the baryon density $\Omega_{b}h^{2}$, and the parameter(s) for the dark component(s), i.e., $\Omega_{c}h^{2}$ for $\Lambda$CDM and $(p_{\rm age},\eta)$ for PUDF. For the neutrino masses, we assume a massive species with minimum mass $0.06\,\mathrm{eV}$ and two massless species. In the context of the $\Lambda$CDM model, we define the matter abundance $\Omega_{m}=\Omega_{b}+\Omega_{c}$ for brevity. Here we do not include $\Omega_{\nu}$ in the definition of $\Omega_{m}$ because we are more interested in matching matter density at high redshift where neutrinos are relativistic.
II Theoretical Comparison between PUDF and $\Lambda$CDM
II.1 PUDF basics
PUDF generalizes the original PAge approximation by adding the radiation and neutrino contribution at high redshift. The Hubble parameter is given by
$H^{2}(z)$ $=$ $H_{0}^{2}\left[\Omega_{\gamma}+\sum_{i=1}^{3}\Omega_{\nu,i}\frac{\mathrm{I}_{\rho}\left(\frac{m_{\nu,i}}{(1+z)T_{\nu}}\right)}{\mathrm{I}_{\rho}\left(\frac{m_{\nu,i}}{T_{\nu}}\right)}\right](1+z)^{4}$ (1)
$+H_{\rm PAge}^{2}(z),$
where $m_{\nu,i}$ is the neutrino mass of the $i$-th species; $T_{\nu}=T_{\rm CMB}\left(\frac{4}{11}\right)^{1/3}\approx 1.95\,\mathrm{K}$ is the effective temperature for neutrino momentum distribution. The neutrino density integral is
$\mathrm{I}_{\rho}(\lambda)\equiv\frac{1}{2\pi^{2}}\int_{0}^{\infty}\frac{x^{2}\sqrt{x^{2}+\lambda^{2}}}{e^{x}+1}~{}\mathrm{d}x.$ (2)
The contribution from baryon and dark fluid is encoded in the $H_{\rm PAge}^{2}(z)$ term. The function $H_{\rm PAge}(z)$ is given by two parameters $(p_{\rm age},\eta)$ and an auxiliary variable $\beta$ running from 00 to $p_{\rm age}$.
$H_{\mathrm{PAge}}$ $=$ $H_{0}\sqrt{1-\Omega_{\nu}-\Omega_{\gamma}}$ (3)
×$\times\left[1+\frac{2}{3}\left(1-\eta\frac{\beta}{p_{\mathrm{age}}}\right)\left(\frac{1}{\beta}-\frac{1}{p_{\mathrm{age}}}\right)\right],$
$1+z$ $=$ $\left(\frac{p_{\rm age}}{\beta}\right)^{2/3}$ (4)
×$\times e^{-\frac{\eta}{3}\left[\left(\frac{\beta}{p_{\rm age}}\right)^{2}-1\right]-\left[p_{\rm age}-\frac{2}{3}(1+\eta)\right]\left(\frac{\beta}{p_{\rm age}}-1\right)}.$
Here the parameter $p_{\rm age}$ is approximately the age of the universe in unit of $H_{0}^{-1}$ and $\eta$ is a phenomenological parameter describing the deviation from the Einstein-de Sitter universe. The running variable $\beta$ is approximately $H_{0}t$.
The density of the unified dark fluid is given by
$\rho_{d}(z)=\frac{3}{8\pi G}H_{\mathrm{PAge}}^{2}-\rho_{b}(z),$ (5)
where is the physical baryon density $\rho_{b}(z)$ is
$\rho_{b}(z)=\rho_{\rm crit}\Omega_{b}(1+z)^{3}\propto\Omega_{b}h^{2}(1+z)^{3}.$ (6)
The pressure of the dark fluid, $p_{d}$, is derived from the continuity equation
$\dot{\rho}_{d}+3H(\rho_{d}+p_{d})=0,$ (7)
and the equation of state for the unified dark fluid is given by
$w\equiv\frac{p_{d}}{\rho_{d}}=\frac{1+z}{3\rho_{d}}\frac{\mathrm{d}\rho_{d}}{\mathrm{d}z}-1.$ (8)
The linear perturbation equations of the unified dark fluid in the synchronous gauge are
$\dot{\delta}$ $=$ $-(1+w)\left(\theta+\frac{\dot{h}_{i}^{i}}{2}\right)-3\frac{\dot{a}}{a}\left(c_{\mathrm{s,eff}}^{2}-w\right)\delta$ (9)
$-9\left(\frac{\dot{a}}{a}\right)^{2}\left(c_{\mathrm{s,eff}}^{2}-c_{\mathrm{s,ad}}^{2}\right)\left(1+w\right)\frac{\theta}{k^{2}},$
$\dot{\theta}$ $=$ $-\frac{\dot{a}}{a}\,\left(1-3c_{\mathrm{s,eff}}^{2}\right)\theta+\frac{c_{\mathrm{s,eff}}^{2}}{1+w}\,k^{2}\delta-k^{2}\sigma\,,$ (10)
where $\delta=\delta\rho_{d}/\rho_{d}$ is the relative density perturbation, $\theta$ is the velocity divergence of the dark fluid, $k$ is the comoving wavenumber, $h_{i}^{i}$ is the trace of the metric perturbations, and $\sigma$ is the shear perturbations of the fluid which is assumed to be negligible in this work. The adiabatic sound speed of the fluid $c_{\mathrm{s,ad}}$ is specified as
$c_{\mathrm{s,ad}}^{2}=\frac{\dot{P}}{\dot{\rho}}=w-\frac{\dot{w}}{3H\left(1+w\right)},$ (11)
The effective sound speed of the unified dark fluid rest frame $c_{\mathrm{s,eff}}^{2}$ is assumed to be zero, too.
II.2 Matching the primary CMB
In the high-redshift limit where $\beta\sim H_{0}t\ll 1$, we may expand Eqs. (3-4) to the linear order of $\beta$ and obtain
$H_{\rm PAge}^{2}$$\approx$ $\frac{4H_{0}^{2}(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}(1+z)^{3}$ (12)
×$\times\left[1+\left(6-\frac{4(1+\eta)}{p_{\rm age}}\right)\beta\right].$
In the pre-recombination epoch where $z\gtrsim 1000$, the $O(\beta)$ correction is below $10^{-4}$ level. Thus, to a very good approximation, $H_{\rm PAge}^{2}$ is proportional to $(1+z)^{3}$ and the unified dark fluid behaves like a CDM component. If we define an effective CDM abundance
$\Omega_{c,\rm eff}=\frac{4(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}-\Omega_{b},$ (13)
the physical density of the dark fluid in the pre-recombination epoch can be written in a familiar way
$\rho_{d}\rvert_{\text{high }z}\approx\rho_{\rm crit}\Omega_{c,\rm eff}(1+z)^{3}.$ (14)
The primary CMB power spectrum relies on the primordial seeds, the pre-recombination physics, the conversion from the physical scale on the last-scattering surface to the observed angular scale, and the scattering between CMB photons and the reionized electrons in the late universe. The parameters controlling these effects are listed in Table 1. It is clear that if we match $\Omega_{c,\rm eff}h^{2}$ in PUDF to $\Omega_{c}h^{2}$ in $\Lambda$CDM, and fix all the other parameters, PUDF and $\Lambda$CDM should predict almost identical primary CMB power spectra with a relative difference less than $O(10^{-4})$. In other words, to match the primary CMB power spectrum to $\Lambda$CDM prediction, $p_{\rm age}$ and $\eta$ should satisfy the constraint
$\left.\frac{4(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM},$ (15)
which simplifies to
$\left.\frac{4}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM},$ (16)
if $\Omega_{\nu}$ and $\Omega_{\gamma}$ are negligible.
Table 1: Parameters controlling primary CMB power spectrum
physical effects parameters
primordial seeds $A_{s}$ and $n_{s}$
pre-recombination physics $\Omega_{b}h^{2}$, $T_{\rm CMB}$, neutrino masses, $\Omega_{c,\rm eff}h^{2}$ for PUDF or $\Omega_{c}h^{2}$ for $\Lambda$CDM
angular scale conversion $\theta_{*}$
reionization $\tau_{\rm re}$
We use Eq. (15) to test the modified Boltzmann code CLASS in Wang et al. [49] and find an $O(10^{-3})$ relative difference between PUDF and $\Lambda$CDM primary CMB power spectra. Further investigation shows that this inconsistency is due to the usage of the subpackage HyRec, which contains a hard-coded $w_{0}w_{a}$ cosmology and therefore can be incompatible with modifications in CLASS. To fix this problem, we replace HyRec with the adapted version of RecFAST in CLASS, which reads cosmology from CLASS. The updated code agrees well with the theoretical expectation that once Eq. (15) is satisfied, the relative difference in primary CMB power spectra of PUDF and $\Lambda$CDM does not exceed $O(10^{-4})$. Figure 1 shows an example where PUDF is matched to the Planck 2018 bestfit $\Lambda$CDM model [1].
Figure 1: Comparison of the primary CMB $TT$ and $EE$ power spectra of PUDF and $\Lambda$CDM when the matching condition (15) is applied. The lower panels give the relative difference ($\Delta C_{\ell}^{\mathrm{TT}}=C^{\mathrm{TT}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{TT}}_{\ell,\Lambda\mathrm{CDM}},\,\Delta C_{\ell}^{\mathrm{EE}}=C^{\mathrm{EE}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{EE}}_{\ell,\Lambda\mathrm{CDM}}$).
II.3 Matching late-universe observables
For a given $\Omega_{m}\rvert_{\rm\Lambda CDM}$, Eq. (16) does not fix $p_{\rm age}$ and $\eta$. We may choose another constraint to match more observables between PUDF and $\Lambda$CDM. For instance, we may match the deceleration parameter $q_{0}=\frac{a\ddot{a}}{\dot{a}^{2}}$ in PUDF and $\Lambda$CDM. In the case of negligible $\Omega_{\nu}$ and $\Omega_{\gamma}$, the $q_{0}$ matching condition is
$\left.\frac{4(1-\eta)}{9p_{\rm age}^{2}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM}.$ (17)
In the original work on PAge where only late universe observables were used, the primary-CMB matching condition (16) was not considered. Instead, the age of the universe in unit of $H_{0}$ was matched [56]. Ignoring the radiation and neutrinos, the age matching condition is
$\left.p_{\rm age}\right|_{\rm PUDF}=\left.\frac{2}{3\sqrt{1-\Omega_{m}}}\ln\frac{1+\sqrt{1-\Omega_{m}}}{\sqrt{\Omega_{m}}}\right|_{\rm\Lambda CDM}.$ (18)
Figure 2: The primary-CMB matching condition (16), $q_{0}$ matching condition (17) and age matching condition (18) for $\Omega_{m}=0.28$ (left panel), $\Omega_{m}=0.315$ (middle panel) and $\Omega_{m}=0.35$ (right panel), respectively.
In Figure 2 we plot the matching conditions for primary CMB, $q_{0}$ and age for a few representative $\Omega_{m}$ values. It is nontrivial to observe that the three conditions almost intersect at one point, where both early- and late-universe observables match well between PUDF and $\Lambda$CDM. It has been shown in Huang [56] that BAO and SN observables can be matched to percent-level accuracy between PAge and $\Lambda$CDM.
While the background evolution is matched between PUDF and $\Lambda$CDM, the abundance and equation of state of the unified dark fluid in PUDF are very different from those of dark matter in $\Lambda$CDM. We may expect very different density perturbations of the dark components in the two models. However, density perturbations of the dark components are not directly observable. What can be observed are the density perturbations of baryonic matter and the bending of the light due to gravitational lensing, both of which track the gravitational potential $\phi$ if anisotropic stress can be ignored. The linear growth of $\phi$ in general depends on the total density perturbation $\delta\rho_{\rm tot}$, the total pressure perturbation $\delta p_{\rm tot}$, and the expansion history of the universe [59]. On sub-horizon scales where the gauge-dependence of $\delta\rho_{\rm tot}$ and $\delta p_{\rm tot}$ can be ignored, we may use the Poisson equation to eliminate the dependence on $\delta\rho_{\rm tot}$ [59]. Thus, in models such as PUDF and $\Lambda$CDM where the rest-frame pressure perturbations are assumed to be negligible, the evolution of $\phi$ on sub-horizon scales only depend on the expansion history of the universe. In other words, for background-matched PUDF and $\Lambda$CDM, the linear growth of gravitational potential is also approximately matched. This has been numerically verified in Wang et al. [49] where the baryon power spectrum in PUDF was shown to be similar to that in $\Lambda$CDM. In Figure 3 we show that the CMB lensing deflection power spectrum in PUDF and $\Lambda$CDM are similar, too.
Figure 3: Comparison of CMB deflection power spectrum $C_{\ell}^{dd}$ of PUDF and $\Lambda$CDM when the primary-CMB matching condition (15) and the $q_{0}$ matching condition (17) are applied. The lower panel gives the relative difference($\Delta C_{\ell}^{\mathrm{dd}}=C^{\mathrm{dd}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{dd}}_{\ell,\Lambda\mathrm{CDM}}$).
III Parameter Inference
In this section we update the parameter inference for PUDF after fixing the minor error in the Boltzmann code. For a fair comparison we use the same combination of CMB + BAO + SN + CC that have been used in Wang et al. [49]. These datasets include Planck TTTEEE and lensing likelihoods [60, 61], the Pantheon+ compilation of Type Ia supernovae [62], and the recent Dark Energy Spectroscopic Instrument Data Release 1 (DESI DR1, Adame et al. [8]). More details can be found in Wang et al. [49].
The results are listed in Table 2. Compared to the results in Wang et al. [49], the updated PUDF parameters all shift towards $\Lambda$CDM parameters. This is because in Wang et al. [49] the incorrect usage of HyRec leads to mismatched primary CMB and hence biased cosmological parameters. The updated results still favor $\Lambda$CDM model, but the relative Bayesian evidence derived with the MCEvidence code [63, 64] is updated to $\ln B_{\rm\Lambda CDM,PUDF}=3.75$, much less significant than the previous (incorrect) result $\ln B_{\rm\Lambda CDM,PUDF}=6.27$ in Wang et al. [49]. With a Bayesian evidence $\Delta\ln B\approx 3.75$, we may say that $\Lambda$CDM is only mildly preferred by the data and PUDF remains to be an interesting option for future investigation.
Table 2: Constraints on parameters with CMB+BAO+SN+CC
parameter $\Lambda$CDM PUDF (this work) PUDF [49]
$100\Omega_{b}h^{2}$ $2.248\pm 0.013$ $2.251\pm 0.014$ $2.253\pm 0.014$
$\Omega_{c}h^{2}$ $0.11856\pm 0.00074$
$100\theta_{*}$ $1.04203\pm 0.00029$ $1.04206\pm 0.00029$ $1.04217\pm 0.00029$
$\ln[10^{10}A_{s}]$ $3.053\pm 0.015$ $3.053\pm 0.015$ $3.054\pm 0.016$
$n_{s}$ $0.9688\pm 0.0035$ $0.9697\pm 0.0039$ $0.9710\pm 0.0041$
$\tau_{\rm re}$ $0.0595\pm 0.0075$ $0.0599\pm 0.0077$ $0.0599\pm 0.0082$
$p_{\rm age}$ $0.9619\pm 0.0073$ $0.9637\pm 0.0076$
$\eta$ $0.428\pm 0.022$ $0.432\pm 0.023$
$H_{0}$ $68.05\pm 0.33$ $68.14\pm 0.61$ $68.26\pm 0.64$
Table 3 shows the $\chi^{2}$ difference between PUDF and $\Lambda$CDM for various combinations of the data sets. When CMB is used, PUDF seems to struggle with twisting its parameters to simultaneously fit early- and late-universe observables as well as $\Lambda$CDM does. In the last column we replace CMB data with a constraint on $\Omega_{b}h^{2}$ from the big bang nucleosynthesis (BBN) model [65, 66, 67, 68], leaving essentially only constraints on the late-universe expansion history. In this case, PUDF fits the data slightly better because it is easier to adjust the background expansion history in PUDF which contains one more degree of freedom than $\Lambda$CDM. In summary, these results indicate that the difference between PUDF and $\Lambda$CDM is statistical significant only when CMB, BAO and SN are combined together. However, it has been shown that DESI BAO, CMB and SN are not very mutually consistent when $\Lambda$CDM is assumed [8, 9]. This leads to some concern that the statistically significant $\Delta\chi^{2}=6.5$ or $\Delta\ln B=3.75$ may be caused by some unknown systematics in the data, if $\Lambda$CDM is indeed the correct model. To test this, we generate mock data by replacing the central value of all observables in BAO, SN, and CC with the theoretical predictions of the Planck 2018 bestfit $\Lambda$CDM model [1]. With the mock data we find the $\chi^{2}$ difference between PUDF and $\Lambda$CDM decrease to $0.34$. This indicates that PUDF and $\Lambda$CDM can be hardly distinguished with the precision of current data, if $\Lambda$CDM is the correct underlying model. The statistically significant difference between $\Lambda$CDM over PUDF ($\Delta\chi^{2}=6.5$ or $\Delta\ln B=3.75$) we have found with the real data may be a rare statistical fluctuation or an evidence that $\Lambda$CDM is not the correct model.
Table 3: $\chi^{2}$ difference between PUDF and $\Lambda$CDM
data sets CMB+BAO+SN+CC CMB+BAO+CC CMB+SN+CC BBN+BAO+SN+CC
$\chi^{2}_{\rm PUDF}-\chi^{2}_{\rm\Lambda CDM}$ $6.5$ $0.58$ $1.1$ $-1.73$
IV Discussion and Conclusions
In this study we show that both the background expansion history and the linear perturbations in the visible sector of the universe can be tuned to be $\Lambda$CDM-like in the PAge-like unified dark fluid model. We derive matching conditions for primary CMB and late-universe observables. Using the primary-CMB matching condition to test the numerical accuracy of the PUDF Boltzmann code, we find a minor error in the code used in Wang et al. [49]. After fixing the numerical error we update the parameter inference and find that $\Lambda$CDM is mildly favored by the current CMB+BAO+SN+CC data, with a $\Delta\chi^{2}\approx 6.5$.
The similarity between PUDF and $\Lambda$CDM is only in the visible part of the universe and at the linear-perturbation level. In the dark sector and on nonlinear scales, PUDF or in general a unified-dark-fluid model can be very different from $\Lambda$CDM. For instance, we are not sure if there can be unified-dark-fluid halos in the low-redshift universe, and if yes, whether their morphology is close to that in $\Lambda$CDM. The fluid description is a phenomenological large-scale approximation of an underlying fundamental theory which we have not yet specified. Given the tantalizing possibility of testing cosmology in the deep nonlinear regime with the future releases of DESI and other cosmological surveys, it would be an interesting direction to construct an underlying theory of PUDF and make predictions on nonlinear scales.
Acknowledgements.
This work is supported by the National key R&D Program of China (Grant No. 2020YFC2201600), the National Natural Science Foundation of China (NSFC) under Grant No. 12073088, the National SKA Program of China No. 2020SKA0110402, and Guangdong Basic and Applied Basic Research Foundation (Grant No.2024A1515012573).